Show that the de Broglie wavelength of a particle in a one dimensional box in the first
excited state is equal to the length of the box.
1
Expert's answer
2020-12-25T14:07:13-0500

Energy of the particle in a one-dimensional box


"E=\frac{n^2\pi^2\hbar^2}{2mL^2}=\frac{n^2\pi^2h^2}{8\pi^2mL^2}=\frac{n^2h^2}{8mL^2}"


Assume that "E=\frac{p^2}{2m}".


"\frac{p^2}{2m}=\frac{n^2h^2}{8mL^2}\to L^2=\frac{n^2h^2}{4p^2}"


For "n=2"


"L=\frac{h}{p}"


The de Broglie wavelength


"\lambda=\frac{h}{p}"


So, "\lambda=L" . Answer